OpenStudy (anonymous):

The equation of circle having a diameter with endpoints (-2, 1) and (6, 7) is (x - 4)² + (y - 3)² = 25 (x - 2)² + (y - 4)² = 25 (x - 2)² + (y - 4)² = 100

OpenStudy (anonymous):

Here's how I do it

OpenStudy (caitlinnr14):

I think you find the midpoint between the two points for the center. find the distance between the two points for the diameter and divide that in half for the radius which equals 5. The center is (2,4).

OpenStudy (anonymous):

Thank you so much guys :D

OpenStudy (mayankdevnani):

 now you have to find equation of circle whose end points are given. REMEMBER:- If end points are given,then use always DIAMETRIC FORM OF A CIRCLE. which is stated as :- \[\large \bf (x-a)(x-c)+(y-b)(y-d)=0\] where (a,b) and (c,d) are end points. Using this :-(your question), end points are :-(-2,1) and (6,7), so using diametric form,we get \[\large \bf (x+2)(x-6)+(y-1)(y-7)=0\] after simplyfying these,you get your answer :)

OpenStudy (mayankdevnani):

this is second way of answering a question :)

OpenStudy (caitlinnr14):

you explained it the best @mayankdevnani

OpenStudy (mayankdevnani):

thank you :)

OpenStudy (mayankdevnani):

understood? @hyesukm52

OpenStudy (anonymous):

so,.. is the answer (x - 2)² + (y - 4)² = 25 ?

OpenStudy (mayankdevnani):

after simplfy this,you will get opeion B as the answer. But if in the options,my answer is given,then what will you do,so use my solution

OpenStudy (mayankdevnani):

*option

OpenStudy (anonymous):

ohokies thank you so muchh xx

OpenStudy (mayankdevnani):

welcome :)